# AI Makes You Smarter, But None The Wiser: The Disconnect Between Performance and Metacognition

### v1

@@ Abstract.
Optimizing human-AI interaction requires users to reflect on their own performance critically. Our study examines whether people using AI to complete tasks can accurately monitor how well they perform. Participants (N = 246) used AI to solve 20 logical problems from the Law School Admission Test. While their task performance improved by three points compared to a norm population, participants overestimated their performance by four points. Interestingly, higher AI literacy was linked to less accurate self-assessment. Participants with more technical knowledge of AI were more confident but less precise in judging their own performance. Using a computational model, we explored individual differences in metacognitive accuracy and found that the Dunning-Kruger effect, usually observed in this task, ceased to exist with AI use. We discuss how AI levels our cognitive and metacognitive performance and consider the consequences of performance overestimation for designing interactive AI systems that enhance cognition.
@@
copyright: rightsretainedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; April 26 – May 1, 2025; Yokohama, Japanbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26 – May 1, 2025, Yokohama, Japanccs: Human-centered computing Human computer interaction (HCI)
@@
[Schematic Figure]Schematic visualization of our main findings. Participants using AI (ChatGPT-4o) were overconfident. They overestimated their performance, yet their performance improved. The Dunning-Kruger effect was reduced when assisted by AI to complete the task.
@@ 1. Introduction
Humans have always used technologies to extend their abilities (Khettab, 2019; Clark, 2008; Alexandre e Castro, 2024). Recent advances have aimed to improve human performance and productivity in a range of contexts (Zulfikar et al., 2024; Hou et al., 2024; Wang et al., 2020; Perera, 2024). While there is evidence for an improvement in human performance with AI (Bansal et al., 2021; Steyvers et al., 2022; Zulfikar et al., 2024), these performance gains are limited by effective Human-AI Interaction (HAI). Fundamental biases, such as overtrust and overreliance, impair performance (Inkpen et al., 2023) up to the point that the interaction decreases overall performance as compared to having no AI at all (Bastani et al., 2024). This raises an important question of whether users can and do monitor their AI-augmented performance.
@@
Human metacognition research investigates our ability to monitor, evaluate, and regulate our own cognitive processes (Fleming, 2024) and, therefore, has been proposed to be essential in interactive generative AI systems (Tankelevitch et al., 2024). Psychological research on metacognition has shown that people typically estimate themselves to be better than average (Brown, 1986), also called the ”better-than-average effect” (see (Zell et al., 2020)).
@@
In the context of AI, people believe AI improves performance (Kloft et al., 2024), that AI predictions outperform professionals (Shekar et al., 2024) and hope AI will improve their lives (Cave and Dihal, 2019). Research on HAI shows evidence of deficiencies in metacognitive monitoring: users are largely unaware of their performance and their performance improvement with AI. When using AI systems, users tend to overestimate its benefits, even when using a sham AI system (Villa et al., 2023; Kloft et al., 2024; Kosch et al., 2023). However, accurate metacognitive monitoring is crucial for optimizing HAI. Inaccurate evaluation of human-AI composite performance (Engelbart, 1962) can lead to an overreliance on the system, resulting in suboptimal outcomes. From a rational perspective (Oulasvirta et al., 2022), optimal interaction with AI requires that users possess a clear understanding of their performance to adjust their behaviour.
@@
Similarly, metacognitive judgments exhibit considerable individual variability (Ackerman and Thompson, 2017; Toplak et al., 2011), which can relate to cognitive performance (Toplak et al., 2011). The Dunning-Kruger Effect (DKE) describes a cognitive bias where individuals with lower ability overestimate their competence while those with higher ability underestimate it (Kruger and Dunning, 1999). For HAI, a DKE would suggest that low performers may not optimize their interaction with AI due to poor metacognitive monitoring. However, one could argue that if AI interaction improves overall cognitive performance by augmenting intellect (Engelbart, 1962), then metacognitive bias and its link to cognitive performance may disappear. To disambiguate these opposing arguments, we must empirically explore metacognitive monitoring, including metacognitive bias, individual metacognitive accuracy, metacognitive sensitivity, and its relation to performance (DKE) in HAI. We thus study whether AI impacts self-assessments of performance (RQ1), if it affects the ability to distinguish between correct and incorrect judgments (RQ2), and if it amplifies or reduces self-assessment bias between low- and high-performing individuals (RQ3).
@@
We investigate metacognitive monitoring and the opposing perspectives on the DKE by measuring how AI affects users’ metacognitive judgments. We designed an experiment where participants used AI to complete logical reasoning tasks from the Law School Admission Test (LSAT). This setting is analogous to those utilized in prior research on metacognitive abilities and the DKE (Jansen et al., 2021). By analyzing participants’ self-assessments before and after AI interaction, we can directly measure the AI’s influence on participants’ metacognition and study its relation to task performance (DKE).
@@
We found that while AI-assistance improves task performance in the LSAT, it also leads to overestimation of users’ performance (low metacognitive accuracy). Yet, we show using a computational model that the DKE is not only smaller but disappears entirely in our sample while being present in a comparable large-scale sample without AI use (Jansen et al., 2021). Technological knowledge and critical appraisal of AI, as measured by the Scale for the assessment of non-experts’ AI literacy (SNAIL) (Laupichler et al., 2023), increased confidence but decreased the accuracy of self-assessment.
@@
To summarize, although AI has the potential to improve performance and level individual biases in metacognition, it carries the risk of inflated performance assessments. Our paper extends our understanding of metacognitive monitoring in HAI by investigating the interplay between metacognition, cognitive performance and AI literacy. While AI had great benefits for performance and levelled metacognitive deficits, it also brought about the bane of overestimating one’s performance. We discuss how to navigate this trade-off and how to improve metacognitive accuracy to empower users to optimize their interactions with AI. Our contributions and research results are:
@@
  • Highlighting a paradox where higher AI literacy relates to less accurate self-assessment, with participants more confident yet less precise in their performance evaluations.
  • Offering design recommendations for interactive AI systems to enhance metacognitive monitoring, empowering users to critically reflect on their performance.
@@ 2.1. Human Metacognition
Metacognitive processes, such as monitoring or evaluating one’s cognitive operations, are key for problem-solving, learning, and behavior optimization (Fleming, 2024). Two main constructs in research on metacognitive judgments are metacognitive accuracy and sensitivity. Metacognitive accuracy refers to how well self-evaluations align with actual performance, enabling better decision-making when limitations are recognized (Fleming, 2024; Colombatto and Fleming, 2023). Metacognitive sensitivity reflects the ability to distinguish between correct and incorrect judgments, often measured by confidence ratings post-decision (Fleming, 2024). Metacognitive accuracy is shaped by metacognitive bias (consistent over/underestimation of abilities) and noise (random fluctuations in self-assessment) (Fiedler et al., 2019). Bias skews evaluations predictably, while noise introduces inconsistency, reducing sensitivity (Fleming, 2024). The DKE appears in the connection between metacognitive accuracy and skill. It suggests that less-skilled individuals overestimate their performance, while highly competent individuals underestimate theirs (Kruger and Dunning, 1999). Debates exist on whether the DKE is a statistical artifact or if it reflects true population trends (Gignac and Zajenkowski, 2020; Gignac, 2024). Jansen et al. (2021), as well as others (e.g., Burson et al. (2006)) addressed these issues, replicating Kruger and Dunning (1999) with over 3000 participants, confirming the existence of the DKE in verbal and logical reasoning tests. Low performers struggled to estimate their accuracy, while high performers slightly underestimated their performance. Therefore, metacognitive abilities are vital for optimizing cognitive processes and relate to individual performance differences in skill, comprising the DKE.
@@ 2.2. Metacognition in Human-AI Interaction
Augmenting human intellect has been central to HCI, as highlighted by Engelbart and others (Engelbart, 1962; Schmidt, 2017) and metacognition has been studied, albeit seldomly quantitatively, at CHI (Tombaugh et al., 1985; Tankelevitch et al., 2024; Quinn and Russell, 1986; Chan et al., 2014; Corbett and Anderson, 2001; Zhang et al., 2022; Foster et al., 2013; Seitlinger and Ley, 2012; Gooch et al., 2016; Loksa et al., 2016; Lu and Yin, 2021). Recent AI advancements have extended this vision, enabling the processing of multi-modal data—sound, visuals, and text—to solve complex problems across fields like medical treatment (Moor et al., 2023), drug discovery (Mak et al., 2023), and climate change (Kaack et al., 2022), as well as in personal contexts (Draxler et al., 2023). Studies have measured AI’s benefits to human cognitive performance, such as Noy and Zhang (2023) finding a 40% increase in students’ essay-writing speed with AI, and Bastani et al. (2024) reporting a 48% improvement in math performance with ChatGPT. However, students using ChatGPT still performed 17% worse in exams compared to a control group (Bastani et al., 2024), indicating that AI can sometimes hinder skill development (Bastani et al., 2024; Rafner et al., 2022; Kobiella et al., 2024). These issues likely stem from suboptimal interfaces that fail to support metacognition (Tankelevitch et al., 2024; Kosch et al., 2023; Kloft et al., 2024). Research shows that users often overestimate their AI-assisted performance and struggle to monitor or plan interactions effectively (Kosch et al., 2023; Kloft et al., 2024; Villa et al., 2023; Bosch et al., 2024). For instance, Zamfirescu-Pereira et al. (2023) found that users have difficulty crafting effective prompts, while Dang et al. (2023) noted challenges in switching between tasks and writing prompts. Furthermore, explanations from AI systems are often uninformative, ignored, or lead to cognitive biases themselves (Eiband et al., 2019; Wang and Yin, 2021; Vasconcelos et al., 2023; Bertrand et al., 2022). AI literacy, which involves understanding AI concepts and evaluating outputs critically, is also essential for effective interaction (Laupichler et al., 2023). However, its influence on metacognitive judgements in AI-assisted decision-making and interaction optimization is unclear. Therefore, a key limitation in human-AI interaction is primarily metacognitive, involving difficulties in planning, monitoring, and evaluating interactions. Theoretically, AI imposes new metacognitive demands on users (Tankelevitch et al., 2024), requiring enhanced monitoring and control. However, empirical research on metacognitive monitoring is scarce.
3. Method
@@
In the following, we motivate and document our methodological choices when conducting the study. The research software, the analysis with all associated measures and data, can be found at [blinded for review].
3.1. Participants
@@
To explore individual differences in cognitive and metacognitive performance, we recruited a larger sample than typical DKE studies, allowing us to detect differences in metacognitive accuracy across high and low performers (Dunning, 2011; Gignac and Szodorai, 2024) [^1] We powered for the smallest effect of interest which is the DKE. For power analysis, we used bootstrapped samples of Jansen et al. (2021) with sample sizes ranging from 80 to 400 to assess the ability to detect the Dunning-Kruger effect through t-tests across quartiles. We computed the proportion of p-values ¡ .05 to determine the optimal sample size for sufficient statistical power (80%). With this, we found that a sample size of 250-300 participants is optimal for reliably detecting differences between the upper and lower quartiles in metacognitive accuracy..
@@
We recruited 274 English-fluent participants located in the USA through Prolific. We included an attention check, requiring participants to read a short description of the study and task. They then answered two multiple-choice questions, one about the topic (logical reasoning) and another regarding which option to choose when solving the problems (the best one). We excluded thirteen (13) participants due to failing the attention check, as well as two (2) due to erroneous responses (e.g., exceeded the number of possible correct answers in estimating performance) and thirteen (13) due to low completion times.
@@
We further analyzed data from 246 participants (identified as female 114; identified as male 130, identified as non-binary 2; Age: $M$ = 39.85, $SD$ = 14.53). When asked to estimate their English fluency, 218 participants reported themselves as native English speakers, 25 as fully fluent, two as conversationally fluent and one as understanding basic English. No participants preferred not to disclose their language proficiency. 14 participants in our sample reported their highest educational degree to be a doctoral degree, 58 a higher tertiary education degree (Master’s level), 95 a lower tertiary education degree (Bachelor’s level), 52 an upper secondary school / high school and 27 a vocational college degree. 13 participants had taken the LSAT before; their performance was slightly lower, $M$ = 12.38, $SD$ = 3.22, compared to those who have not taken it, $M$ = 13.01, $SD$ = 2.85, thus they were not excluded from the sample. We collected informed consent from each participant before the study in accordance with the Declaration of Helsinki guidelines (Association et al., 2013). Each participant was compensated ¿anonymized¡ upon completion.
3.2. Materials

### v2

# Performance and Metacognition Disconnect when Reasoning in Human-AI Interaction

@@ Abstract.
Optimizing human-AI interaction requires users to reflect on their own performance critically. Our paper examines whether people using AI to complete tasks can accurately monitor how well they perform. In Study 1, participants (N = 246) used AI to solve 20 logical problems from the Law School Admission Test. While their task performance improved by three points compared to a norm population, participants overestimated their performance by four points. Interestingly, higher AI literacy was linked to less accurate self-assessment. Participants with more technical knowledge of AI were more confident but less precise in judging their own performance. Using a computational model, we explored individual differences in metacognitive accuracy and found that the Dunning-Kruger effect, usually observed in this task, ceased to exist with AI. Study 2 (N = 452) replicates these findings. We discuss how AI levels metacognitive performance and consider consequences of performance overestimation for interactive AI systems enhancing cognition.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; April 26 – May 1, 2025; Yokohama, Japanbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26 – May 1, 2025, Yokohama, Japanccs: Human-centered computing Human computer interaction (HCI)
@@
[Schematic Figure]Schematic visualization of our main findings. An interface in the middle with symbols around. Participants using AI (ChatGPT-4o) were overconfident. They overestimated their performance, yet their performance improved. The Dunning-Kruger effect was reduced when assisted by AI to complete the task.
@@ 1. Introduction
Humans have always used technologies to extend their abilities (Khettab, 2019; Clark, 2008; Alexandre e Castro, 2024). Recent advances have aimed to improve human performance and productivity in a range of contexts (Zulfikar et al., 2024; Hou et al., 2024; Wang et al., 2020; Perera, 2024). While there is evidence for an improvement in human performance with AI (Bansal et al., 2021; Steyvers et al., 2022; Zulfikar et al., 2024), these performance gains are limited by effective Human-AI Interaction (HAI). Fundamental biases, such as overtrust and overreliance, impair performance (Inkpen et al., 2023) up to the point that the interaction decreases overall performance as compared to having no AI at all (Bastani et al., 2024; Vaccaro et al., 2024). This raises an important question of whether users can and do monitor their AI-augmented performance.
@@
Human metacognition research investigates the ability to monitor, evaluate, and regulate our own cognitive processes (Fleming, 2024) and, therefore, has been proposed to be essential in interactive generative AI systems (Tankelevitch et al., 2024). Psychological research on metacognition has shown that people typically estimate themselves to be better than average (Brown, 1986), also called the ”better-than-average effect” (see (Zell et al., 2020)).
@@
In the context of AI, people believe AI improves performance (Kloft et al., 2024), that AI predictions outperform professionals (Shekar et al., 2024) and hope AI will improve their lives (Cave and Dihal, 2019). Research on HAI shows evidence of deficiencies in metacognitive monitoring: users are largely unaware of their performance and their performance improvement with AI. Concretely, when using AI systems, users tend to overestimate their benefits, even when using a sham AI system (Villa et al., 2023; Kloft et al., 2024; Kosch et al., 2023). However, accurate metacognitive monitoring is crucial for optimizing HAI. Inaccurate evaluation of human-AI composite performance (Engelbart, 1962) can lead to an overreliance on the system, resulting in suboptimal outcomes. From a rational perspective (Oulasvirta et al., 2022), optimal interaction with AI requires that users possess a clear understanding of their performance to adjust their behavior.
@@
Similarly, metacognitive judgments exhibit considerable individual variability (Ackerman and Thompson, 2017; Toplak et al., 2011), which can relate to cognitive performance (Toplak et al., 2011). The Dunning-Kruger Effect (DKE) describes a cognitive bias where individuals with lower ability overestimate their competence while those with higher ability underestimate it (Kruger and Dunning, 1999). For HAI, a DKE would suggest that low performers may not optimize their interaction with AI due to poor metacognitive monitoring. However, one could argue that if AI interaction improves overall cognitive performance by augmenting intellect (Engelbart, 1962), then metacognitive bias and its link to cognitive performance may disappear. To disambiguate these opposing arguments, we must empirically explore metacognitive monitoring, including metacognitive bias, individual metacognitive accuracy, metacognitive sensitivity, and its relation to performance (DKE) in HAI. We thus conceptually replicate Jansen et al. (2021) in interaction with AI to explore whether AI impacts self-assessments of performance (RQ1: Does interaction with AI affect metacognitive accuracy?), if it affects the ability to distinguish between correct and incorrect judgments (RQ2: Does interaction with AI affect metacognitive sensitivity?), and if it amplifies or reduces self-assessment bias between low- and high-performing individuals (RQ3: Does interaction with AI affect the DKE pattern?).
@@
We investigate metacognitive monitoring and the opposing perspectives on the DKE by measuring how AI affects users’ metacognitive judgments. We designed an experiment where participants used AI to complete logical reasoning tasks from the Law School Admission Test (LSAT). This setting is analogous to those utilized in prior research on metacognitive abilities and the DKE (Jansen et al., 2021). By analyzing participants’ self-assessments after AI interaction, we can directly measure the AI’s influence on participants’ metacognition and study its relation to task performance (DKE).
@@
In Study 1, we found that while using AI improves task performance in the LSAT, it also leads to a large overestimation of users’ performance (low metacognitive accuracy). Yet, we show using a computational model that the DKE is not only smaller but disappears entirely in our sample while being present in a comparable large-scale sample without AI use (Jansen et al., 2021). Technological knowledge and critical appraisal of AI, as measured by the ”Scale for the assessment of non-experts AI literacy” (SNAIL) (Laupichler et al., 2023), increased confidence but decreased the accuracy of self-assessment. In Study 2 (N = 452), where we incentivize metacognitive monitoring with monetary benefits and collect our own non-AI baseline group, we replicate the pattern of results of Study 1.
@@
To summarize, although AI has the potential to improve performance and level individual biases in metacognition, it carries the risk of inflated performance self-assessments. Our paper extends our understanding of metacognitive monitoring in HAI by investigating the interplay between metacognition, cognitive performance, and AI literacy. While AI had great benefits for performance and leveled metacognitive deficits, it also brought about the bane of overestimating one’s performance. We discuss how to navigate this trade-off and how to improve metacognitive accuracy to empower users to make better decisions when using interactive AI. Our contributions and research results are:
@@
  • Highlighting a paradox where higher AI literacy relates to less accurate self-assessment, with participants being more confident yet less precise in their performance evaluations.
  • Offering design recommendations for interactive AI systems to enhance metacognitive monitoring by empowering users to critically reflect on their performance.
@@ 2.1. Human Metacognition
Metacognitive processes (e.g., monitoring or evaluating one’s cognition) are crucial for problem-solving, learning, and behavior optimization (Fleming, 2024). Metacognitive judgments primarily involve accuracy and sensitivity: accuracy measures how closely self-evaluations match actual performance, enabling better decisions when recognizing limitations (Fleming, 2024; Colombatto and Fleming, 2023), while sensitivity assesses the ability to distinguish between correct and incorrect judgments, often via confidence ratings (Fleming, 2024).
Metacognitive accuracy is shaped by bias (systematic over/underestimation) and noise (random fluctuations) (Fiedler et al., 2019), where bias skews evaluations and noise reduces sensitivity (Fleming, 2024). The Dunning–Kruger effect (DKE) links metacognitive accuracy with skill level: low-performing individuals overestimate their performance, while high performing individuals underestimate theirs (Kruger and Dunning, 1999). Although some argue the DKE may be a statistical artefact (Gignac and Zajenkowski, 2020; Gignac, 2024), studies like Jansen et al. (2021) and Ehrlinger et al. (2008) replicated these findings with large samples, confirming the DKE in verbal and logical reasoning tasks.
Although metacognition and the DKE are frequently examined in educational contexts (e.g., see Yang Hansen et al. (2024) in math education and Mahdavi (2014) for an overview), our approach to studying the DKE shifts the focus to a task-specific context. Following Dunning (2011), we replicate the method of Jansen et al. (2021) by concentrating on task performance and ratings of absolute performance estimates after task completion.
In sum, metacognitive abilities are central to optimizing cognition and explain performance differences underlying the DKE.
@@ 2.2. Metacognition in Human-AI Interaction
Augmenting human intellect has been central to HCI, as highlighted by Engelbart and others (Engelbart, 1962; Schmidt, 2017) and metacognition has been studied, albeit seldom quantitatively, at CHI (Tombaugh et al., 1985; Tankelevitch et al., 2024; Quinn and Russell, 1986; Chan et al., 2014; Corbett and Anderson, 2001; Zhang et al., 2022; Foster et al., 2013; Seitlinger and Ley, 2012; Gooch et al., 2016; Loksa et al., 2016; Lu and Yin, 2021).
Augmenting human cognitive processes through AI has recently received attention in CHI. Augmentation has been explored in writing (Reza et al., 2024; Göldi et al., 2024; Li et al., 2024; Lee et al., 2024), reading comprehension (Xiao et al., 2023), reasoning (Danry et al., 2023) and solving math arithmetic tasks (Xu and Zhang, 2023). In addition, research has addressed the psychological challenges in augmenting humans with AI, such as appropriate reliance on AI (He et al., 2023), trust in AI and its impact on performance (Ahn et al., 2024), and inefficiencies in task delegation to AI systems (Pinski et al., 2023). However, there is a notable lack of consideration for metacognition (Tankelevitch et al., 2024), which refers to how people monitor their own performance when augmented with AI and what brings optimal performance in human-AI interaction.
A recent survey by Vaccaro et al. (2024) distinguishes between human-AI synergy – where combined performance surpasses either humans or AI alone – and human-AI augmentation, where humans aided by AI do better than unassisted humans. They found that when humans already outperform AI, adding AI improves the team’s overall performance. However, as AI becomes more powerful, the average performance of these teams declines. Thus, a central challenge in human-AI interaction is achieving synergy when AI models surpass human capabilities.
For LLM’s, studies have measured AI’s benefits to human cognitive performance, such as Noy and Zhang (2023) finding a 40% increase in students’ essay-writing speed with AI, and Bastani et al. (2024) reporting a 48% improvement in math performance with ChatGPT. However, students using ChatGPT still performed 17% worse in exams compared to a control group (Bastani et al., 2024), indicating that AI can sometimes hinder skill development (Bastani et al., 2024; Rafner et al., 2022; Kobiella et al., 2024).
@@
These issues likely stem from suboptimal interfaces that fail to support metacognition (Tankelevitch et al., 2024; Kosch et al., 2023; Kloft et al., 2024). Research shows that users often overestimate their AI-assisted performance and struggle to monitor or plan interactions effectively (Kosch et al., 2023; Kloft et al., 2024; Villa et al., 2023; Bosch et al., 2024). For instance, Zamfirescu-Pereira et al. (2023) found that users have difficulty crafting effective prompts, while Dang et al. (2023) noted challenges in switching between tasks and writing prompts. Furthermore, explanations from AI systems are often uninformative, ignored, or lead to cognitive biases themselves (Eiband et al., 2019; Wang and Yin, 2021; Vasconcelos et al., 2023; Bertrand et al., 2022). AI literacy, which involves understanding AI concepts and evaluating outputs critically, is also essential for effective interaction (Laupichler et al., 2023). However, its influence on metacognitive judgements in AI-assisted decision-making and interaction optimization is unclear. Therefore, a key limitation in human-AI interaction is primarily metacognitive, involving difficulties in planning, monitoring, and evaluating interactions. Theoretically, AI imposes new metacognitive demands on users (Tankelevitch et al., 2024), requiring enhanced monitoring and control. However, empirical research on metacognitive monitoring is scarce.
3. Study 1: Metacognition in Human-AI interaction
We have conducted two studies. One compares a group of participants using AI to Jansen et al. (2021) data on the LSAT (that have not used AI); the other study, Study 2, replicates and extends Study 1.
3.1. Method
@@
In the following, we motivate and document our methodological choices when conducting Study 1. The research software, the analysis with all associated measures and data, can be found at [blinded for review]. Note that both the data and the material of Jansen et al. (2021) are openly available under https://osf.io/er9ms/, which allowed us to closely follow their task environment and sample characteristics for the purpose of Study 1. All data collected for the purpose of our paper and analysis scripts can be found at https://osf.io/svax9/?view_only=cd176c08c2d4422baabf6bf815018bd7.
3.1.1. Participants
@@
To explore individual differences in cognitive and metacognitive performance, we recruited a larger sample than typical DKE studies, allowing us to detect differences in metacognitive accuracy across high and low performers (Dunning, 2011; Gignac and Szodorai, 2024). We powered for the smallest effect of interest, which is the DKE. For power analysis, we used bootstrapped samples of Jansen et al. (2021) with sample sizes ranging from 80 to 400 to assess the ability to detect the DKE through t-tests across quartiles. We computed the proportion of p-values ¡ .05 to determine the optimal sample size for sufficient statistical power (80%). With this, we found that a sample size of 250-300 participants is optimal for reliably detecting differences between the upper and lower quartiles in metacognitive accuracy..
@@
We recruited 274 English-fluent participants located in the USA through Prolific. We included an attention check, requiring participants to read a short description of the study and task. They then answered two multiple-choice questions, one about the topic (logical reasoning) and another regarding which option to choose when solving the problems (the best one). We excluded thirteen (13) participants due to failing the attention check, as well as two (2) due to erroneous responses (e.g., exceeding the number of possible correct answers in estimating performance) and thirteen (13) due to low completion times.
@@
We further analyzed data from 246 participants (identified as female 114; identified as male 130, identified as non-binary 2; Age: $M$ = 39.85, $SD$ = 14.53). When asked to estimate their English fluency, 218 participants reported themselves as native English speakers, 25 as fully fluent, two as conversationally fluent and one as understanding basic English. No participants preferred not to disclose their language proficiency. 14 participants in our sample reported their highest educational degree to be a doctoral degree, 58 a higher tertiary education degree (Master’s level), 95 a lower tertiary education degree (Bachelor’s level), 52 an upper secondary school / high school and 27 a vocational college degree. 13 participants had taken the LSAT before; their performance was slightly lower, $M$ = 12.38, $SD$ = 3.22, compared to those who have not taken it, $M$ = 13.01, $SD$ = 2.85, thus they were not excluded from the sample. We collected informed consent from each participant before the study in accordance with the Declaration of Helsinki guidelines (Association et al., 2013). Each participant was compensated 6.5 pounds per hour. In accordance with the [Blinded] national guidelines (the [Blinded] National Board on Research Integrity), this study did not require ethics approval as it involved minimal risk to participants, with no intervention beyond standard practice and no collection of sensitive personal data.
3.1.2. Materials